# Some arithmetic aspects of ortho-integral surfaces

### v1

@@ 5.2. Verifying arithmeticity of the double of OI pants and OI one-holed torus
To compute the arithmetic invariants of the genus-two surface obtained by doubling either a pair of pants $P$ or a one-holed torus $T$, we need the reflections across the boundary components of $P$ and $T$. To that end, we consider the following matrices
@@
For $*\in\{\alpha,\beta,\gamma\}$, we can check that $r_{*}$ fixes ${\bf v}_{*}$ and ${\bf w}_{*}$ and defines reflections across the sides of $H(a,b,c)$ opposite to *, see Figure 2.
The description of OI one-holed tori and OI pants in terms of an orthobasis makes it natural to describe the corresponding Fuchsian groups using the quadratic form $q=q_{(a,b,c)}$.
@@
  • Proof.
    The proof is a direct computation verifying the three conditions in 5.3. Let $P$ be an OI pair of pants and $T$ be an OI one-holed torus. The Fuchsian group that corresponds to the double of a pair of pants $P$ and a one-holed torus $T$ is $\Delta_{P}$ and $\Delta_{T}$ given by 5.4. Using Equation 24, Equation 23 and the classification of OI pants and OI one-holed tori, we verify that the trace set of $\Delta_{P}$ and $\Delta_{T}$ is integral for any OI pair of pants and OI one-holed torus. Similarly, we verify that the invariant trace field of $\Delta_{P}$ and $\Delta_{T}$ are all $\mathbb{Q}$ which totally real. The last condition in 5.3 is vacuous. ∎
The following tables record the arithmetic data of OI pants and one-holed torus. The first column records the cosh-length of an ortho-basis for the surfaces. The second column records a Hilbert symbol of the invariant quaternion algebra of the surface computed using Equation 26. The third column records the finite primes in $\mathbb{Z}$ over which $A$ ramifies, i.e. the completion of $A$ at $p$ is a division algebra. For a fixed number field, the ramification set is a complete invariant of the invariant quaternion algebra [MR03]. The final column records the trace set of the genus 2 surface group that corresponds to doubling $P$ or $T$.
Table 2. Classification of OI pants with associated invariants
@@ 5.3. Constructing infinitely many commensurability classes of OI surfaces
  • Proof.
    By construction, the universal cover $\widetilde{X}$ of $X$ admits a tiling by an all right-angled hyperbolic hexagon $H(a,a,a)$ for $a\in\{2,3\}$. Thus, $\widetilde{X}$ can be identified with a convex subset of $P(\mathcal{C}^{+})$ tilled by $H(a,a,a)$ for $a\in\{2,3\}$. The boundary of $\widetilde{X}$ consists of geodesics in $\mathcal{U}^{\perp}$. The set of cosh-lengths of orthogeodesics in $X$ is precisely the set of cosh-distances among geodesics in $\partial\widetilde{X}$. By 5.6, these cosh-distances are one of the coordinates of a vector in $\mathcal{U}=\Lambda\cdot\{\vec{\bf u}_{\alpha},\vec{\bf u}_{\beta},\vec{\bf u}_{\gamma}\}$. For $a\in\{2,3\}$, the group $\Lambda$ has integral entries. That is, $\Lambda\leq\operatorname{O}^{+}(q,\mathbb{Z})$. Therefore, $\mathcal{U}\subseteq\mathbb{Z}^{3}$. It follows that $X$ is an ortho-integral surface. ∎
@@
Figure 3. Gluing copies of OI pants $P(a,a,a)$ without twist, for a fixed $a\in\{2,3\}$.
@@
Finally, as an application of 5.2 and 5.6, we construct examples of $d$-AOI but not OI pants using quadratic form.
@@
  • Proof.
    In the case of $P(\frac{n+1}{n},\frac{n+1}{n},\frac{n+1}{n})$, the group $G_{P}\leq\operatorname{O}^{+}(q;\mathbb{Z})$ where $q$ is the quadratic form given in Equation 17. Furthermore, one can check that $G_{P}$ preserves the $\mathbb{Z}$-lattice $(\frac{1}{n}\mathbb{Z})^{3}$. This implies that the $G_{P}$-orbits of $\{\vec{\bf u}_{\alpha},\vec{\bf u}_{\beta},\vec{\bf u}_{\gamma}\}$ are contained in $(\frac{1}{n}\mathbb{Z})^{3}$. Since the entries of the vectors in the $G_{P}$-orbits of $\{\vec{\bf u}_{\alpha},\vec{\bf u}_{\beta},\vec{\bf u}_{\gamma}\}$ consist of the cosh-length spectrum of the orthogeodesics of $P(\frac{n+1}{n},\frac{n+1}{n},\frac{n+1}{n})$. These pants provide examples of AOI surfaces but not OI surfaces. The argument is similar in the case of $P(\frac{2n+3}{2n+1},\frac{2n+3}{2n+1},\frac{2n+3}{2n+1})$. ∎
@@
Using the same argument as in 5.7, we can glue copies of the pair of pants $P(\frac{3}{2},\frac{3}{2},\frac{3}{2})$ together without any twist along their boundaries, either by gluing a copy to itself or to another, to obtain additional 2-AOI surfaces with higher topological complexity.

### v2

@@ 5.2. Verifying arithmeticity of the double of OI pair of pants and OI one-holed torus
Here, the invariant trace field $k\Gamma$ is $k\Gamma=\mathbb{Q}(\{\operatorname{tr}_{\operatorname{SL}}\gamma^{2}\mid\gamma\in\Gamma\})$. The invariant quaternion algebra $A\Gamma$ is defined to be $A\Gamma=\{\sum a_{i}\gamma_{i}\mid a_{i}\in k\Gamma,\gamma_{i}\in\Gamma\}$. To compute the arithmetic invariants of the genus-two surface obtained by doubling either a pair of pants $P$ or a one-holed torus $T$, we need the reflections across the boundary components of $P$ and $T$. To that end, we consider the following matrices
@@
For $\in\{\alpha,\beta,\gamma\}$, we can check that $r_{*}$ fixes $\bf v$$_{*}$ and $\bf w$$_{*}$ and defines reflections across the sides of $H(a,b,c)$ opposite to *, see Figure 3.
The description of OI one-holed tori and OI pairs of pants in terms of an orthobasis makes it natural to describe the corresponding Fuchsian groups using the quadratic form $q=q_{(a,b,c)}$.
@@
The proof is a direct computation verifying the three conditions in 5.3. Let $P$ be an OI pair of pants and $T$ be an OI one-holed torus. The Fuchsian group that corresponds to the double of $P$ and the double of $T$ is $\Delta_{P}$ and $\Delta_{T}$ given by 5.4. Using Equation 24, Equation 23 and the classification of OI pairs of pants and OI one-holed tori, we verify that the trace set of $\Delta_{P}$ and $\Delta_{T}$ is integral. After doubling $P$ or $T$, every orthogeodesics in $P$ and in $T$ becomes a closed geodesic in the genus two surface, the trace of these particular geodesics is twice the cosh of the length of the orthogeodesic from which they originate. To check that the trace set of $\Delta_{P}$ and of $\Delta_{T}$ is integral, we use [MR03, Lemma 3.5.2] which produces a finite set of traces that we need to check for integrality. In particular, for a 4-generated group $\Gamma=\langle w,x,y,z\rangle$, it suffices to check integrality for the traces of the elements
$\{w,x,y,z,wx,wy,wz,xy,xz,yz,wxy,wxz,wyz,xyz,wxyz\}.$
In all cases, the traces of all elements in $\Delta_{P}$ or $\Delta_{T}$ are all algebraic integers. See LABEL:tab:OIPantsInvariants and LABEL:tab:OIToriInvariants for the resulting computations.
To compute the invariant trace field, [MR03, Lemma 3.5.3] and [MR03, Lemma 3.5.5] imply that for a non-elementary subgroup $\Gamma=\langle w,x,y,z\rangle\leq\operatorname{PSL}_{2}(\mathbb{C})$, it suffices to compute the traces of
$\{w^{2},x^{2},y^{2},z^{2},w^{2}x^{2},w^{2}y^{2},w^{2}z^{2},x^{2}y^{2},x^{2}z^{2},y^{2}z^{2},w^{2}x^{2}y^{2},w^{2}x^{2}z^{2},w^{2}y^{2}z^{2},x^{2}y^{2}z^{2}\}.$
The resulting computations show that the invariant trace field of $\Delta_{P}$ and $\Delta_{T}$ are all $\mathbb{Q}$ which are totally real. The last condition in 5.3 is vacuous. ∎
The following tables record the arithmetic data of OI pair of pants and one-holed torus. The first column records the cosh-length of an ortho-basis for the surfaces. The second column records a Hilbert symbol of the invariant quaternion algebra of the surface computed using Equation 26. The third column records the finite primes in $\mathbb{Z}$ over which $A$ ramifies, i.e. the completion of $A$ at $p$ is a division algebra. For a fixed number field, the ramification set is a complete invariant of the invariant quaternion algebra [MR03, Theorem 2.7.5] and therefore uniquely specifies the algebra up to isomorphism. The final column records the trace set of the genus 2 surface group that corresponds to doubling $P$ or $T$.
Table 2. Classification of OI pairs of pants with associated invariants
@@ 5.3. Constructing infinitely many commensurability classes of OI surfaces
By construction, the universal cover $\widetilde{X}$ of $X$ admits a tiling by an all right-angled hyperbolic hexagon $H(a,a,a)$ for $a\in\{2,3\}$. Thus, $\widetilde{X}$ can be identified with a convex subset of $P(\mathcal{C}^{+})$ tilled by $H(a,a,a)$ for $a\in\{2,3\}$. The boundary of $\widetilde{X}$ consists of geodesics in $\mathcal{U}^{\perp}$. The set of cosh-lengths of orthogeodesics in $X$ is precisely the set of cosh-distances among geodesics in $\partial\widetilde{X}$. By 5.6, these cosh-distances are one of the coordinates of a vector in $\mathcal{U}=\Lambda\cdot\{{\bf u}_{\alpha},{\bf u}_{\beta},{\bf u}_{\gamma}\}$. For $a\in\{2,3\}$, the group $\Lambda$ has integral entries. That is, $\Lambda\leq\operatorname{O}^{+}(q,\mathbb{Z})$. Therefore, $\mathcal{U}\subseteq\mathbb{Z}^{3}$. It follows that $X$ is an ortho-integral surface. ∎
@@
Figure 4. Gluing copies of OI pair of pants $P(a,a,a)$ without twist, for a fixed $a\in\{2,3\}$.
@@
The sequence of hyperbolic surfaces $\{X_{n}\}$, defined as in the paragraph prior to Figure 4, consists of all OI surfaces by 5.7. The universal covers of these surfaces $X_{n}$ are pairwise non-isometric by 5.11. It follows that the surfaces $\{X_{n}\}$ are pairwise non-commensurable. ∎
@@
Remark 5.13.
Thanks to the referee’s suggestion, another way to show the sequence of OI surfaces $\{X_{n}\}$ are pairwise incommensurable is by comparing the orthokissing number and area:
$\operatorname{okiss}(X_{n})=3\cdot 2^{n-1},\qquad\operatorname{area}(X_{n})=2\pi(3\cdot 2^{n}-2),$
for $n\geq 1$. If $X_{n}$ and $X_{m}$ are commensurable then the ratio $\operatorname{area}/\operatorname{okiss}$ coincides, hence
$\frac{3\cdot 2^{n}-2}{3\cdot 2^{n-1}}=\frac{3\cdot 2^{m}-2}{3\cdot 2^{m-1}}\ \Rightarrow\ n=m.$
Thus, the sequence $\{X_{n}\}$ are pairwise incommensurable.
Finally, as an application of 5.2 and 5.6, we construct examples of pairs of pants that are $d$-AOI but not OI using quadratic forms.
@@
In the case of $P(\frac{n+1}{n},\frac{n+1}{n},\frac{n+1}{n})$, the group $G_{P}\leq\operatorname{O}^{+}(q;\mathbb{Z})$ where $q$ is the quadratic form given in Equation 17. Furthermore, one can check that $G_{P}$ preserves the $\mathbb{Z}$-lattice $(\frac{1}{n}\mathbb{Z})^{3}$. This implies that the $G_{P}$-orbits of $\{{\bf u}_{\alpha},{\bf u}_{\beta},{\bf u}_{\gamma}\}$ are contained in $(\frac{1}{n}\mathbb{Z})^{3}$. Since the entries of the vectors in the $G_{P}$-orbits of $\{{\bf u}_{\alpha},{\bf u}_{\beta},{\bf u}_{\gamma}\}$ consist of the cosh-length spectrum of the orthogeodesics of $P(\frac{n+1}{n},\frac{n+1}{n},\frac{n+1}{n})$. These pairs of pants provide examples of AOI surfaces but not OI surfaces. The argument is similar in the case of $P(\frac{2n+3}{2n+1},\frac{2n+3}{2n+1},\frac{2n+3}{2n+1})$. ∎
@@
Using the same argument as in 5.7, we can glue copies of the pair of pants $P(\frac{3}{2},\frac{3}{2},\frac{3}{2})$ together without any twists along their boundaries, either by gluing a copy to itself or to another, to obtain additional 2-AOI surfaces with higher topological complexity.